OpenStudy (anonymous):

Calculus: rate of change Two sides of a triangle have lengths 12m and 12m. The angle between them is increasing at a rate of 2 degrees per minute. How fast is the length of the third side increasing when the angle between the sides of fixed length is 60 deg.?

OpenStudy (amistre64):

law of cosines should be useful for this

OpenStudy (amistre64):

\[c^2=a^2+b^2-2ab~\cos(\alpha)\] \[2cc'=2aa'+2bb'-2ab'~\cos(\alpha)-2a'b~\cos(\alpha)-2ab\alpha'~\sin(\alpha)\] \[cc'=aa'+bb'-ab'~\cos(\alpha)-a'b~\cos(\alpha)-ab\alpha'~\sin(\alpha)\] since a' and b' are not changing, they equal 0 \[cc'=-ab\alpha'~\sin(\alpha)\] \[c'=-\frac{ab}{c}\alpha'~\sin(\alpha)\]

OpenStudy (amistre64):

i think i got a bad negative during that lol

OpenStudy (amistre64):

a,b,c = 12; alpha=60, alpha'=2 \[c'=\frac{12(12)}{12}2~\sin(60)\] \[c'=24\sin(60)\]

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

step by step checking it out.........

OpenStudy (anonymous):

c' =ab/c sine x....yes, but where do you get c=12?

OpenStudy (amistre64):

when the angle is 60 degrees, you have an equilateral triangle ...

OpenStudy (amistre64):



OpenStudy (anonymous):

Oh my goodness........I knew that! Thanks